OpenStudy (anonymous):

Please help me evaluate this integral. √3−4/x−√+6/e(2x)−5/x−12x3+1/x2−x+4)

OpenStudy (anonymous):

Here is the integral

OpenStudy (anonymous):

\[\sqrt[3]{?x^2}-4/\sqrt{x}+6/e^(2x)-5/x-12x^3+1/x^2-x+4)\]

OpenStudy (anonymous):

you can basically rewrite everything in the integrand with positive/negative exponents then just use the power rule....

OpenStudy (anonymous):

this is just the integrand: \(\large \sqrt[3]{x^2}-\frac{4}{\sqrt x}+\frac{6}{e^{2x}}-\frac{5}{x}-12x^3+\frac{1}{x^2}-x+4\) =\(\large x^{\frac{2}{3}}-4x^{-\frac{1}{2}} \) .... taking too long... you know what i mean....

OpenStudy (anonymous):

@dpalnc thanks for you help